The tangent of p is -1/radical 3 and the secant of p is -2/ radical 3. What is the sine of p?
Answer by stanbon(48558)

The tangent of p is -1/radical 3 and the secant of p is -2/ radical 3. What is the sine of p?
---------------------
tan(p) = -1/sqrt(3) = y/x
sec(p) = -2/sqrt(3) = r/x implies r = 2 and x = -sqrt(3)
y = sqrt(r^2-x^2) = sqrt[4 - 3] = 1
----------------------------------
If tan is negative and sec is neg, p must be in the 2nd quadrant
where sin is positive.
sin(P) = y/r = 1/2
===============================
Cheers,
Stan H.